Two balls are drawn in succession without replacement from a box containing 9 black balls (B) and 7 white balls (W). Let Z be the random variable representing the number of white balls. Possible Outcomes Value of the Random Variable Z Number of White Balls (Z) Probab


1
Expert's answer
2022-05-06T12:59:21-0400

Possible outcomes of the Random Variable Z are: "0,1,2."


"P(Z=0)=P(BB)"

"=\dfrac{9}{16}(\dfrac{8}{15})=\dfrac{3}{10}"

"P(Z=1)=P(BW)+P(WB)"

"=\dfrac{9}{16}(\dfrac{7}{15})+\dfrac{7}{16}(\dfrac{9}{15})=\dfrac{21}{40}"



"P(Z=2)=P(WW)"

"=\dfrac{7}{16}(\dfrac{6}{15})=\dfrac{7}{40}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 z & 0 & 1 & 2 \\ \hline
 p(z) & 3/10 & 21/40 & 7/40 \\
\end{array}"